A.construct a table of values for y=x^2-4x+3. use 0 B. Using your table from A, draw a table a graph of y=x^2-4x+3 in the space at the right.
c. based oon your graph and table, when x^2-4x+3=0 what is x?__________________
D. what is the formula of the line of reflection?_____________
E. what are the coordinates of the turning point?_____________
please please please answer my questions.

Answer by drk(1908)

(A) First we construct a table of values as follows using y=x^2-4x+3
x: . . . . . -2 . . . . . -1 . . . . .0 . . . . .1 . . . . 2 . . . . . .3
y: . . . . .. 14 . . . . .8 . . . . . . 3 . . . . 0 . . . . -1 . . . . . .0
(B) you can graph this next part based on this table.
(C) When Y = 0, then x = 3 or x = 1
(D) Line of reflection (axis of symmetry) is at X = 2
(E) If you mean max or min, then the vertex is a minimum at (2,-1)